OpenStudy (anonymous):

Evaluate \(f(5)\) if \[f(x)-x\,f(-x)=x^2\]

OpenStudy (michele_laino):

Please note thatif I make this substitution: x--->-x, I can re-write your equation as below: \[f(-x)+xf(x)=x ^{2}\] now I multiply both sides by x, and I get: \[xf(-x)+x ^{2}f(x)=x ^{3}\] adding, side by side, that last above equation and your equation, I get this equation: \[(x ^{2}+1)f(x)=x ^{3}+x ^{2}\] and: \[f(x)=\frac{ x ^{3}+x ^{2} }{ x ^{2}+1 }\] so substituting x=5, I get: \[f(5)=\frac{ 125+25 }{ 25+1 }=\frac{ 150 }{ 26 }=\frac{ 75 }{ 13 }\]

OpenStudy (anonymous):

That's correct! Nice work

OpenStudy (michele_laino):

Thanks! @SithsAndGiggles

OpenStudy (anonymous):

I saw this listed in a math contest from a while back and it looked like an interesting problem :)

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

nice question! :)